OpenStudy (anonymous):

What is the solution of the equation 6x − 3 = -51?

OpenStudy (anonymous):

A. -9 B. -8 C. 8 D. 9

OpenStudy (anonymous):

add 3 to both sides 6x= -48 divide by 6 on both sides x = -8

OpenStudy (anonymous):

Solve the equation 2(3x + 7) = -22

OpenStudy (anonymous):

distribute that 2 in front of 3x +7 6x + 14 = -22 subtract 14 on both sides 6x = -36 divide by 6 on both sides x = -6

OpenStudy (anonymous):

What is the equation of a line passing through (-3, 7) and having a slope of -1/5?

OpenStudy (anonymous):

the standard form is:\[y-y_1 = m(x-x_1)\]where m is your slope and y_1 and x_1 are your points

OpenStudy (anonymous):

A. y = -5x + 22 B. y = 5x + 22 C. y= 1/5x + 32/5 D. y= -1/5x + 32/5 E. y= -1/5x + 22

OpenStudy (anonymous):

just put in your points and slope into the formula i gave you

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Okay I got the answer thanks though! But here's another problem.

OpenStudy (anonymous):

whats the question?

OpenStudy (anonymous):

What is the equation of this line? A. y= =1/3x - 3/4 B. y= -1/3x - 4 C. y= 1/3x + 4 D.y = 3x – 4

OpenStudy (anonymous):

\[slope = \frac{ y_2-y_1 }{ x_2-x_1 }\] find the slope using the 2 points it gives you, then use the standard form formula again using the slope and whatever point you want

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Got it!